# Limits on Dark Matter Compact Objects implied by Supermagnified Stars in Lensing Clusters

### v1

@@ Abstract
Supermagnified stars are gravitationally lensed individual stars that are located close to a caustic of a lensing galaxy cluster, and have their flux magnified by a large enough factor (typically $\sim$ 1000) to make them detectable with present telescopes. The maximum magnification is limited by microlensing caused by intracluster stars or other compact objects, which create a network of corrugated critical lines with an angular width proportional to the surface density of microlenses. We consider a set of 9 cases of supermagnified stars reported in the literature, and derive an upper limit on the surface density of compact objects, such as primordial black holes, that might be present as a fraction of the dark matter in addition to known intracluster stars. Any such additional compact objects would widen the corrugated critical line network and therefore the width of the distribution of supermagnified stars around the modeled critical lines of the lens. We find that any compact objects, including primordial black holes, with masses above $\sim 10^{-6}\,M_{\odot}$ (for which the microcaustics are smaller than the typical angular size of supermagnified stars) cannot account for more than $\sim$ 3% of the dark matter.
@@ 1 Introduction
Dark matter may be in the form of compact objects, which would be detectable through gravitational lensing. The baryon density derived from primordial nucleosynthesis and CMB observations rules out compact objects made of baryonic matter, but primordial black holes (PBH) formed in the early universe from the collapse of the hot plasma in regions of large amplitude primordial perturbations can account for cold dark matter without requiring the addition of any particles to the Standard Model of particle physics (Hawking, 1971). Although a number of astrophysical observations have constrained the contribution of PBH to the dark matter of the Universe (see Carr et al., 2021, for a review), the asteroid-mass window $10^{-16}\,M_{\odot}<m_{b}<10^{-10}\,M_{\odot}$ remains completely open. Higher mass PBH are ruled out as constituents of all the dark matter by microlensing surveys in the Magellanic Clouds, although a small fraction of $\sim$ 10% of dark matter in compact objects is still allowed (Alcock et al., 1998, 2001; Tisserand et al., 2007; Wyrzykowski et al., 2011a, b; Griest et al., 2014; Smyth et al., 2020).
@@
In this paper we analyze the most interesting cases of highly magnified stars that have been reported in the literature to derive an upper limit to the fraction of dark matter in lensing clusters that may be in the form of dark, compact objects. Section 2 reviews the properties of the network of micro-critical lines that are relevant to derive this upper limit, and introduces an analytical approximation for the density of supermagnified images around this network. Our method to derive an upper limit on compact objects is described in Section 3, together with the list of supermagnified stars we use in this work. Our results are presented in Section 4, and we summarize our conclusion in Section 6.
@@ 2 The distribution of highly magnified images around the smooth critical line
We now discuss the expected distribution of highly magnified images in the observed image plane of a lens. If we imagine first a smooth lens, without any small-scale structure, highly magnified images are located in the vicinity of the critical lines (or lines where the determinant of the magnification matrix is zero), which are smooth curves. A real lensing cluster has some fraction of its mass in compact objects: at the very least, intracluster stars are present. The macro-model of the lens is defined to have the surface density resulting from smoothing the point masses over a scale that is much smaller than the total lens deflection but much larger than the separation among neighboring microlenses.
We follow Venumadhav et al. (2017), hereafter V17, to establish our nomenclature and notation. The macro-critical line (MaCL) is defined as the critical line of the lens macro-model. The true critical line takes into account the individual lensing stars in the intracluster light without smoothing, and is called the micro-critical line (MiCL). The MiCL is highly complex, with a corrugated band of angular width $r_{w}$ where it meanders around many of the point masses. This angular width is given by (see eq. 22 in V17)
@@
where $\kappa_{\star}$ is the contribution to the lensing convergence from the smoothed point masses, and $g$ is the angular gradient modulus of the eigenvalue of the macro-model magnification matrix that vanishes at the MaCL (note that $g$ is written as $d$ in V17). In the region within a separation $r<r_{w}$ from the MaCL, the MiCLs are connected to each other and form an intricate, corrugated network that is affected everywhere by many point masses. Outside the corrugated network ($r>r_{w}$), the MiCLs are usually disconnected and form closed curves around each point mass, with the shape of the infinite symbol in the approximation of a single point mass with fixed external shear (Chang & Refsdal, 1979).
@@
We first review the analytical derivation for how the maximum magnification reached in MiCL crossings depends on $r/r_{w}$. The maximum magnification $\mu_{p}$ that is reached when a source star crosses a MiCL depends on the local eigenvalue gradient $g_{\star}$, computed including the point masses with no smoothing (whereas $g$ is the eigenvalue gradient of the macro-lens model). We assume all microlenses have the same mass, corresponding to an Einstein radius $\theta_{\star}$. The peak magnification is $\mu_{p}\propto(g_{\star}\sin\alpha_{\star})^{-1/2}$ (eq. 27 in V17), where $\alpha_{\star}$ is the angle formed by the MiCL and the principal axis of the eigenvalue that cancels on the MiCL. The average separation between neighboring point masses is $\theta_{\star}/\kappa_{\star}^{1/2}$. Within the corrugated band, the MiCL is a curve meandering around point masses at this typical separation, where the typical shear from point masses is $\lambda_{\star}\sim\kappa_{\star}$ and its gradient is $g_{\star}\sim\kappa_{\star}^{3/2}/\theta_{\star}$. Therefore, the distribution of peak magnifications of MiCL crossings is roughly constant within the width $r_{w}$, with the typical value scaling as $\mu_{p}\propto\theta_{\star}^{1/2}\kappa_{\star}^{-3/4}$. We assume that the distribution of the angle $\alpha_{\star}$ is random and independent of $r$, so it does not affect the proportionality relations we are discussing.
@@
However, when moving outside the network at a separation $r>r_{w}$ from the MaCL, MiCLs are closed curves around individual point masses of a characteristic size $\theta_{i}\sim\theta_{\star}(\kappa_{\star}r/r_{w})^{-1/2}$, where the characteristic shear is now $\lambda_{\star}\sim\kappa_{\star}(r/r_{w})$ and the eigenvalue gradient is typically $g_{\star}\sim\lambda_{\star}/\theta_{i}\sim(\kappa_{\star}\,r/r_{w})^{3/2}/\theta_{\star}$. The reason is that the shear of a point mass drops with the angular separation as $\theta_{i}^{-2}$, and the shear from microlenses that needs to be added to the macro-model to reach a zero of the magnification eigenvalue increases as $r/r_{w}$. The peak magnification that is reached is therefore $\mu_{p}\propto\theta_{\star}^{1/2}(\kappa_{\star}\,r/r_{w})^{-3/4}$. When expressed as a function of the source separation from the macroscopic caustic, $y\propto r^{2}$, the peak magnification drops as $\mu_{p}\propto y^{-3/8}$ outside the corrugated band of micro-caustics, and is flat inside (see eq. 27 in V17).
@@
Nevertheless, what we actually want to infer here is not the dependence of peak magnifications on $r$, but the distribution of highly magnified images around the MaCL that reach a magnification above some threshold value, $\mu>\mu_{t}$. For this purpose, we compute first the fraction of the image plane area where the magnification is above $\mu_{t}$, in the limit of high $\mu_{t}$ when this fraction is small, as a function of $r$. The region around each MiCL that has magnification $\mu>\mu_{t}$ has a width proportional to $g_{\star}^{-1}$, assuming the source star is small enough to be fully inside the corresponding width around the microcaustic in the source plane. Within $r<r_{w}$ this fraction is roughly constant because the distribution of $g_{\star}$ is also roughly independent of $r$. Outside the corrugated band, the typical value of $g_{\star}$ on MiCLs grows as $(r/r_{w})^{3/2}$, and the length of MiCLs is proportional to their typical size around each point mass, or to $(r/r_{w})^{-1/2}$. Therefore, the fraction of the area with $\mu>\mu_{t}$, scales as $g_{\star}^{-1}(r/r_{w})^{-1/2}\sim(r/r_{w})^{-2}$.
@@
Because we have computed this area fraction at fixed $\mu$, the density of micro-images above $\mu_{t}$ from sources that are randomly distributed in the source plane must also scale in the same way: roughly constant within $r_{w}$ and falling as $(r/r_{w})^{-2}$ at $r>r_{w}$. Based on these approximations, we adopt the following model for computing likelihood functions of the observed positions of highly magnified images:
$p_{i}(r)\,dr=\frac{2}{\pi}\,\frac{dr/r_{w}}{1+(r/r_{w})^{2}}~{}.$ (2)
@@
We note that this expression for the function $p_{i}(r)$ assumes that the source star is small enough not to have reached the maximum possible magnification, which is produced when the stellar disk lies on the micro-caustic. Once this maximum magnification is reached, the probability density $p_{i}$ drops much faster with $r$ than in equation (2), because most MiCL become too small to reach the required magnification.
@@ 3 Corrugated band widths
We consider these events are most likely to be true supermagnified stars. Supermagnified stars will follow the distribution of equation (2) only when their detection requires a minimum magnification threshold to make them visible, and the source size is small enough to remain unresolved when magnified (otherwise the images will be considered as typical surface brightness irregularities that are seen in many lensed galaxies). Single star candidates may be confused with other objects, mainly compact globular clusters that remain unresolved, but this is unlikely for most of these candidates since clusters that are compact enough are rare (see Welch et al., 2022a). The main criteria to consider these events as good supermagnified star candidates are photometric data that are consistent with the spectra of some of the most luminous stars (generally supergiants or very massive main-sequence stars), and variability on timescales of $\sim$ days that is consistent with being caused by MiCL crossings of sources with radii typical of the most luminous known stars, with a lack of observed counterimages which are typically not seen because of insufficient magnification.
These candidates generally satisfy these criteria. There are doubts, however, on the nature of the two Spock objects and Godzilla (events labeled as 2, 3 and 6 in Table 1), which may be more consistent with flaring luminous blue variables (similar to Eta Carinae) or recurrent novae (Rodney et al., 2018; Diego et al., 2022), especially in view of their high brightness. In addition, for these cases it is difficult to determine precisely where the MaCL lies (see in particular the discussion in Diego et al., 2022), and therefore the separation $r$ from the observed variable images to the MaCL. The position of the MaCL can often be determined rather precisely when pairs of images of surface brightness irregularities can be identified, and the MaCL must lie in the middle of such pairs when the pairs are much closer together compared to the size of the MaCL. This generally gives a high confidence in the modelled separations $r$ listed in column 9 of Table 1 for the remaining 9 cases, all of which have consistent photometry and variability to be microlensing events of single stars or compact systems of few stars (binaries or multiple hierarchical binaries).
@@
We list also the convergence contributed by stars that is implied by the intracluster light brightness near the image, $\kappa_{\star}$, that is reported in the listed references for 6 of the 9 events. In most cases, values of the surface mass density in microlenses ($\Sigma_{\star}$) is reported after assuming a Salpeter or Chabrier initial mass function, and we then compute $\kappa_{\star}=\Sigma_{\star}/\Sigma_{\text{crit}}$. This measurement of $\Sigma_{\star}$ is not reported yet for the other 3 cases.
@@
The predicted corrugated band half widths, $r_{w}$, are then computed using equation (1), and are shown as black horizontal bars in Figure 1, for all the 9 events we use. For event number 6 we did not have the magnification matrix of the lens model available and the red bar indicates the mean of the other lenses.
We also show as red horizontal bars the same bandwith $r_{w}$ when we assume that compact objects account for 1% of the total mass surface density of the lensing model (i.e., $\kappa_{\star}=0.01\kappa_{0}$). The observed angular separations between the star images and each model MaCL are shown as blue dots.
We see that generally, the images are always observed at the expected separations from the MaCL, compared to $r_{w}$. There are two cases where the image is substantially further, events 6 and 11. As discussed earlier, we do not consider event 6 in view of the modeling uncertainties; in fact the large separation from the MaCL for this event is quite unreliable and may actually be much smaller, as discussed by Diego et al. (2022). However, we think that event 11, and also event 9 with a more moderate separation outside the corrugated band, are real cases of microlensed stars that can occur further from the MaCL according to equation (2), as we further discuss below.
@@
Figure 1: Angular distance $r$ from each of the 12 highly magnified star images to the model macro critical line (MaCL) (blue dots), compared to the half-width of the corrugated band of micro-critical lines ($r_{w}$, black bars) when the surface density $\kappa_{\star}$ is that corresponding to the intracluster light, and the same corrugated band half-width for $\kappa_{\star}=0.01\ \kappa_{0}$ (red bars). Events 2, 3 and 6 are excluded from our analysis because of uncertainties on the nature of the sources and the lens model of the MaCL.

### v2

@@ Abstract
Supermagnified stars are gravitationally lensed individual stars that are located close to a caustic of a lensing galaxy cluster, and have their flux magnified by a large enough factor (typically $\sim$ 1000) to make them detectable with present telescopes. The maximum magnification is limited by microlensing caused by intracluster stars or other compact objects, which create a network of corrugated critical lines with an angular width proportional to the surface density of microlenses. We consider a set of 9 cases of supermagnified stars reported in the literature, and derive an upper limit on the surface density of compact objects, such as primordial black holes, that might be present as a fraction of the dark matter in addition to known intracluster stars. Any such additional compact objects would widen the corrugated critical line network and therefore the width of the distribution of supermagnified stars around the modeled critical lines of the lens. We find that any compact objects, including primordial black holes, with masses above $\sim 10^{-6}\,M_{\odot}$ (below which the microcaustics are closer together than the typical angular size of supermagnified stars) cannot account for more than $\sim$ 2% of the dark matter.
@@ 1 Introduction
Dark matter may be in the form of compact objects, which would be detectable through gravitational lensing. The baryon density derived from primordial nucleosynthesis and CMB observations rules out compact objects made of baryonic matter, but primordial black holes (PBH) formed in the early universe from the collapse of the hot plasma in regions of large amplitude primordial perturbations can account for cold dark matter without requiring the addition of any particles to the Standard Model of particle physics (Hawking, 1971). Although a number of astrophysical observations have constrained the contribution of PBH to the dark matter of the Universe (see Carr et al., 2021, for a review), the asteroid-mass window $10^{-16}\,M_{\odot}<m_{b}<10^{-10}\,M_{\odot}$ remains completely open. Higher mass PBH are ruled out as constituents of all the dark matter by microlensing surveys in the Magellanic Clouds, although a small fraction of $\sim$ 10% of dark matter in compact objects is still allowed (Alcock et al., 1998, 2001; Tisserand et al., 2007; Wyrzykowski et al., 2011a, b; Griest et al., 2014; Smyth et al., 2020). This fraction is even lower in the stellar mass range from the most recent results of the OGLE experiment (Mróz et al., 2024a, b).
@@
In this paper we analyze the most interesting cases of highly magnified stars that have been reported in the literature to derive an upper limit to the fraction of dark matter in lensing clusters that may be in the form of dark, compact objects. Section 2 reviews the properties of the network of micro-critical lines that are relevant to derive this upper limit, and introduces an analytical approximation for the density of supermagnified images around this network. Our method to derive an upper limit on compact objects is described in Section 3, together with the list of supermagnified stars we use in this work. Our results are presented in Section 4, with a discussion on the dependence on the lens mass in Section 5, and we summarize our conclusion in Section 6.
@@ 2 The distribution of highly magnified images around the smooth critical line
We now discuss the expected distribution of highly magnified images in the observed image plane of a lens.
In general, highly magnified images are located in the vicinity of the critical lines, defined as the curves in the image plane where the determinant of the magnification matrix is zero. If the lens surface density is a differentiable function, without structure at scales less than some characteristic smoothing scale, the critical lines are also smooth and differentiable curves. A real lensing cluster, however, has some fraction of its mass in compact objects: at the very least, intracluster stars acting as point masses are present. It is then convenient to define the macro-model of the lens as a smoothed version of the true lens, with a surface density resulting from smoothing the point masses over a scale that is much smaller than the total lens deflection, but much larger than the separation among neighboring microlenses.
We follow Venumadhav et al. (2017), hereafter V17, to establish our nomenclature and notation. The macro-critical line (MaCL) is defined as the critical line of the lens macro-model, including the smoothed contribution from point masses, which we refer to as microlenses hereafter. The smooth critical line (SCL), on the other hand, is the critical line of all the rest of the lens mass, without including the microlenses, which is assumed to be intrinsically smooth on small scales. The true critical line takes into account the microlenses (the known intracluster stars, plus any other hypothetical compact objects that may account for some fraction of the dark matter), and is called the micro-critical line (MiCL). The MiCL is highly complex, and forms a corrugated band of microcaustics that meander around many microlenses. As discussed in V17, this corrugated band is centered on the SCL and not the MaCL, with an angular width given by (see eq. 22 in V17)
@@
where $\kappa_{\star}$ is the contribution to the lensing convergence from the smoothed point masses, and $g$ is the angular gradient modulus of the eigenvalue of the macro-model magnification matrix that vanishes at the MaCL (note that $g$ is written as $d$ in V17). Defining $s$ to be the angular separation from the SCL in the lens plane, with a positive sign for points outside the SCL and negative sign inside, the MiCLs are typically connected to each other to form an intricate, corrugated network in the region $|s|<r_{w}$, which is affected everywhere by many point masses. Outside the corrugated network, at $|s|>r_{w}$, the MiCLs are usually disconnected and form closed curves around each point mass, with the shape of the infinite symbol in the approximation of a single point mass with fixed external shear (Chang & Refsdal, 1979). The total width of the corrugated band is therefore $2r_{w}$. Because the macro-model has a greater surface density by $\kappa_{\star}$ than the smooth component, the MaCL is external to the SCL and is shifted outwards by an angular separation $r_{w}$ from the SCL, just at the outer edge of the corrugated band.
@@
We first review the analytical derivation for how the maximum magnification reached in MiCL crossings depends on $s/r_{w}$. The maximum magnification $\mu_{p}$ that is reached when a source star of angular radius $\theta_{s}$ crosses a MiCL depends on the local eigenvalue gradient $g_{\star}$, computed including the point masses with no smoothing (whereas $g$ is the eigenvalue gradient of the macro-lens model). We assume for simplicity that all microlenses have the same mass, corresponding to an Einstein radius $\theta_{\star}$ (we discuss in Section 5 the likely effect of a distribution of microlens masses). The peak magnification is $\mu_{p}\propto(g_{\star}\theta_{s}\sin\alpha_{\star})^{-1/2}$ (eq. 27 in V17), where $\alpha_{\star}$ is the angle formed by the MiCL and the principal axis of the eigenvalue that cancels on the MiCL. The average separation between neighboring point masses is $\theta_{\star}/\kappa_{\star}^{1/2}$. Within the corrugated band, the MiCL is a curve meandering around point masses at this typical separation, where the typical shear from point masses is $\lambda_{\star}\sim\kappa_{\star}$ and its gradient is $g_{\star}\sim\kappa_{\star}^{3/2}/\theta_{\star}$. Therefore, the distribution of peak magnifications of MiCL crossings is roughly constant within the width $r_{w}$, with the typical value scaling as
$\mu_{p}\propto(\theta_{\star}/\theta_{s})^{1/2}\kappa_{\star}^{-3/4}\qquad(|s|<r_{w})~.$ (2)
We assume that the distribution of the angle $\alpha_{\star}$ is random and independent of $r$, so it does not affect the proportionality relations discussed here.
@@
However, when moving outside the network at a separation $|s|>r_{w}$ from the SCL, MiCLs are closed curves around individual point masses of a characteristic size $\theta_{i}\sim\theta_{\star}(\kappa_{\star}|s|/r_{w})^{-1/2}$, where the characteristic shear is now $\lambda_{\star}\sim\kappa_{\star}(|s|/r_{w})$ and the eigenvalue gradient is typically $g_{\star}\sim\lambda_{\star}/\theta_{i}\sim(\kappa_{\star}\,|s|/r_{w})^{3/2}/\theta_{\star}$. The reason is that the shear of a point mass drops with the angular separation as $\theta_{i}^{-2}$, and the shear from microlenses that needs to be added to the macro-model to reach a zero of the magnification eigenvalue increases as $s/r_{w}$. The peak magnification that is reached is therefore reduced to
$\mu_{p}\propto(\theta_{\star}/\theta_{s})^{1/2}(\kappa_{\star}\,|s|/r_{w})^{-3/4}\qquad(|s|>r_{w})~.$ (3)
When expressed as a function of the source separation from the caustic of the smooth model, $y\propto s^{2}$, the peak magnification drops as $\mu_{p}\propto y^{-3/8}$ outside the corrugated band of micro-caustics, and is flat inside (see eq. 27 in V17).
@@
Nevertheless, what we actually want to infer here is not the dependence of peak magnifications on $s$, but the distribution of highly magnified images around the SCL that reach a magnification above some threshold value, $\mu>\mu_{t}$. For this purpose, we compute first the fraction of the image plane area where the magnification is above $\mu_{t}$, in the limit of high $\mu_{t}$ when this fraction is small, as a function of $s$. The region around each MiCL that has magnification $\mu>\mu_{t}$ has a width proportional to $g_{\star}^{-1}$, assuming the source star is small enough to be fully inside this width from the microcaustic in the source plane. Within $|s|<r_{w}$ this fraction is roughly constant because the distribution of $g_{\star}$ is also roughly independent of $s$. Outside the corrugated band, the typical value of $g_{\star}$ on MiCLs grows as $(|s|/r_{w})^{3/2}$, and the length of MiCLs is proportional to their typical size around each point mass, or to $(|s|/r_{w})^{-1/2}$. Therefore, the fraction of the area with $\mu>\mu_{t}$, scales as
${\rm Area}\sim g_{\star}^{-1}(|s|/r_{w})^{-1/2}\sim(|s|/r_{w})^{-2}~.$ (4)
@@
Because we have computed this area fraction at fixed $\mu$, the density of micro-images above $\mu_{t}$ from sources that are randomly distributed in the source plane must also scale in the same way: roughly constant in $|s|<r_{w}$ and falling as $(|s|/r_{w})^{-2}$ at $|s|>r_{w}$. Based on these approximations, we adopt the following model for computing likelihood functions of the observed positions of highly magnified images:
$p_{i}(s)\,ds=\frac{1}{\pi}\,\frac{ds/r_{w}}{1+(s/r_{w})^{2}}~.$ (5)
@@
We note that this expression for the function $p_{i}(s)$ assumes that the source star is small enough not to have reached the maximum possible magnification, which is produced when the stellar disk lies on the micro-caustic. Once this maximum magnification is reached, the probability density $p_{i}$ drops much faster with $s$ than in equation (5), because most MiCL become too small to reach the required magnification.
Our adoption of equation (5) for the functional form of the probability distribution of magnified images is done for simplicity in this brief and necessarily preliminary study. A more thorough analysis will naturally benefit from a detailed functional form derived from numerical simulations of microlensing.
Observations usually provide good constraints on the position of the MaCL of lensing models, when pairs of images of the same extended object are observed on large scales compared to the microlenses, which are at equal separations $r$ from the MaCL. As discussed earlier, the separation from the SCL is $s=r+r_{w}$, where $s$ and $r$ are defined to be positive outside the SCL and MaCL respectively, and negative inside. Therefore, our model for the distribution of magnified images expressed in terms of the separation from the MaCL is
$p_{i}(r)\,dr=\frac{1}{\pi}\,\frac{dr/r_{w}}{1+(1+r/r_{w})^{2}}~.$ (6)
@@ 3 Corrugated band widths
We consider these events are most likely to be true supermagnified stars. Supermagnified stars should follow the distribution approximated by equation (6) only when their detection requires a minimum magnification threshold to make them visible, and the star size is smaller than the separation from the microcaustic when magnified. Single supermagnified star candidates may be confused with other objects, mainly compact globular clusters that remain unresolved, but this is unlikely for most of these candidates because clusters that are compact enough are rare (see Welch et al., 2022). The main criteria to consider these events as good supermagnified star candidates are photometric data that are consistent with the spectra of some of the most luminous stars (generally supergiants or the most massive main-sequence stars), and variability on timescales of $\sim$ days that is consistent with being caused by MiCL crossings of sources with radii typical of these most luminous stars, with a lack of observed counterimages which are typically not seen owing to insufficient magnification.
The candidates in Table 1 generally satisfy these criteria. There are doubts, however, on the nature of the two Spock objects and Godzilla (events labeled as 2, 3 and 6 in Table 1), which may be more consistent with flaring luminous blue variables (similar to Eta Carinae) or recurrent novae (Rodney et al., 2018; Diego et al., 2022), especially in view of their high brightness. In addition, for these cases it is difficult to determine precisely where the MaCL lies (see in particular the discussion in Diego et al., 2022), and therefore the separation $r$ from the observed variable images to the MaCL. The position of the MaCL is best determined when pairs of images of surface brightness irregularities in the source galaxy can be identified, which indicate that the MaCL lies precisely in between the image pair (assuming the pair separation is much less than the MaCL size, so that a linear variation of the magnification matrix from the MaCL is a good approximation). This generally gives a high confidence in the modelled separations $r$ listed in column 9 of Table 1 for the remaining 9 cases, all of which have consistent photometry and variability to be microlensing events of single stars or compact systems of few stars (binaries or multiple hierarchical binaries).
From the discussion in the previous section, we note that the separation from the SCL is obtained from the separation from the MaCL as $s=r+r_{w}=r+\kappa_{\star}/g$, and that the sign of $r$ listed in Table 1 is positive outside the MaCL. The corrugated band where the magnifications are highest and the largest density of supermagnified stars are expected is centered on the SCL, at $s=0$.
@@
We list also the convergence contributed by stars that is implied by the intracluster light brightness near the image, $\kappa_{\star}$, that is reported in the listed references for 5 of the 9 considered events. Here we simply list the surface density values directly as reported by the authors in each reference where it is given, although surface densities are obviously inferred from the observed surface brightness, generally assuming a Chabrier initial mass function, or a Salpeter one with a calibrated mass lower limit. This should be improved in the future using the same standardized initial mass function for all cases. We then compute $\kappa_{\star}=\Sigma_{\star}/\Sigma_{\text{crit}}$. This measurement of $\Sigma_{\star}$ is not reported yet for the other 4 cases.
@@
The predicted corrugated band half widths, $r_{w}$, are then computed using equation (1), and are indicated at the position $r=-2r_{w}$ as black horizontal bars in Figure 1, for the 5 events with a known value of $\kappa_{\star}$. In this way, the dashed line at $r=0$ ($s=+r_{w}$) indicates the outer edge of the corrugated band, and the bars at $s=r+r_{w}=-r_{w}$ indicate the inner edge, while the band center is in between at $r=-r_{w}$. We also show as red horizontal bars the bandwith corresponding to a total surface density of compact objects equal to 1% of the total mass surface density of the lensing model (i.e., $\kappa_{\star}=0.01\kappa_{0}$), again at $r=s-r_{w}=-2r_{w}$. The observed angular separations between the star images and each model MaCL are shown as blue dots. For event number 6 we did not have the magnification matrix of the lens model available and the red bar indicates the mean of the other lenses.
We see that generally, the images are observed at locations compatible with the expected center and width of the corrugated band. For the distribution assumed in equation (6), we expect half of the images to lie between the dashed line at $r=0$ and the black or red bars depending on our assumed $\kappa_{\star}$. As mentioned in Section 2, the nature of event 6 is questionable, and the only other event that is a substantial outlier is event 11, which would lie very far from the corrugated band if we believe the very small value of $\kappa_{\star}$ contributed by intracluster stars given by Meena et al. (2023a). Event 11 may be a true outlier in the distribution of highly magnified stars or may have its stellar surface density underestimated.
From this list of cases, it is not yet possible to check if most magnified stars are inside the MaCL, at $r<0$, as expected, but this should become clear as more cases of supermagnified stars with good models and well measured $\kappa_{\star}$ become available. If we take the red bars as an example, the assumption of $\kappa_{\star}=0.01\kappa_{0}$ results in 5 out of 9 cases within the band, with 3 outside the outer edge and 1 inside the inner edge, compared to the prediction of 50% of images in the corrugated band from the simple model of equation (6).
@@
Figure 1: Angular separation $r$ (blue dots) from each of the 12 highly magnified star images to the model macro critical line (MaCL, grey dashed line, which is the outer edge of the corrugated band at $r=0$), compared to the inner edge of the corrugated band of micro-critical lines (black bars) when the surface density $\kappa_{\star}$ is that of the reported intracluster light, and when it is $\kappa_{\star}=0.01\ \kappa_{0}$ (red bars). The corrugated band is centered at $r=-r_{w}$ (the SCL), while its inner edge shown by the bars is at $r=-2r_{w}$. Events 2, 3 and 6 are excluded from our analysis because of uncertainties on the nature of the sources and the lens model of the MaCL.